# Quantum Channel Testing in Average-Case Distance

### v1

@@ 2.3 Query models for channel testers
A (deterministic) ancilla-free, non-adaptive channel tester making $n$ queries to a channel $\mathcal{M}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ is a tuple consisting of the following:
@@
A (deterministic) ancilla-assisted, incoherent, non-adaptive channel tester making $n$ queries to a channel $\mathcal{M}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ is a tuple consisting of the following:
@@
A (deterministic) ancilla-assisted, coherent, non-adaptive channel tester making $n$ queries to a channel $\mathcal{M}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ is a tuple consisting of the following:
@@
A (deterministic) ancilla-assisted, coherent, adaptive channel tester making $n$ queries to a channel $\mathcal{M}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ is a tuple consisting of the following:
@@ 2.4 Von Neumann entropy
We will use von Neumann entropy to prove our results about tomography.
@@
$\mathopen{}\mathclose{{}\left|S\mathopen{}\mathclose{{}\left(\rho^{\prime}}\right)-S\mathopen{}\mathclose{{}\left(\sigma^{\prime}}\right)}\right|\leq x\log\mathopen{}\mathclose{{}\left(d^{\prime}-1}\right)+1\leq x\log\mathopen{}\mathclose{{}\left(d^{\prime}}\right)+1.$
@@ 3 Lower bounds for channel certification in diamond distance
It is straightforward to verify that $\boldsymbol{\mathcal{M}}$ is completely positive and trace-preserving. One may alternatively verify that $\boldsymbol{\mathcal{M}}$ is a channel by interpreting it as the following sequence of physical operations: with probability $1-\varepsilon$ apply $\mathcal{N}$, and with probability $\varepsilon$ apply the channel that first performs the PVM $(\boldsymbol{\phi},I-\boldsymbol{\phi})$ on the input state, and then outputs $\boldsymbol{\psi}$ if the measurement outcome was $\boldsymbol{\phi}$ and outputs $\mathcal{N}$ applied to the post-measurement state if the measurement outcome was $I-\boldsymbol{\phi}$. Thus $\boldsymbol{\mathcal{M}}$ mirrors the behavior of $\mathcal{N}$ except in the neighborhood of $\boldsymbol{\phi}$.
@@
$\geq 2\varepsilon(1-1/d_{\mathrm{out}})$
$\geq\varepsilon.$
@@
where the probability is over both the choice of $\boldsymbol{\mathcal{M}}$ and the randomness of the output measurement. Recalling 2.6, write $T=(\mathcal{V}_{1},\dotsb,\mathcal{V}_{n},P)$ where $n$ is the number of queries made by $T$; our goal is to prove that $n\geq\Omega(\sqrt{d}_{\mathrm{in}}/\varepsilon)$.
@@
$\leq\varepsilon\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\mathcal{N}\mathopen{}\mathclose{{}\left((\boldsymbol{\phi}\otimes I)\xi(\boldsymbol{\phi}\otimes I)-\xi(\boldsymbol{\phi}\otimes I)-(\boldsymbol{\phi}\otimes I)\xi}\right)}\right\|_{1}+\operatorname*{tr}\mathopen{}\mathclose{{}\left((\boldsymbol{\phi}\otimes I)\xi}\right)}\right]$ triangle inequality
@@ 4.2 Relation to average-case distance between unitaries
A survey by Montanaro and de Wolf [MdW13, Section 5.1.1], drawing on results of Low [Low2009Avg] and Wang [Wang2011Unitary], defines the quantity
$D(U,V)\coloneqq\sqrt{1-\frac{1}{d^{2}}\mathopen{}\mathclose{{}\left|\operatorname*{tr}\mathopen{}\mathclose{{}\left(U^{\dagger}V}\right)}\right|^{2}}$
as the natural notion of distance between unitaries $U,V\in\mathbb{C}^{d\times d}$ in the context of quantum property testing. These works collectively proved the equivalent characterizations
$D(U,V)=\frac{1}{\sqrt{2d}}{\|U\otimes U^{\dagger}-V\otimes V^{\dagger}\|}_{2}=\sqrt{\frac{d+1}{4d}\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|U\boldsymbol{\psi}U^{\dagger}-V\boldsymbol{\psi}V^{\dagger}}\right\|_{1}^{2}}\right]}$ (20)
where $\boldsymbol{\psi}\in\mathsf{D}(d)$ is Haar random, of which the latter justifies $D$ as an “average-case distance”. ACID distance generalizes $D$ because if channels $\mathcal{U},\mathcal{V}\in\mathsf{C}(d,d)$ conjugate by unitaries $U,V\in\mathbb{C}^{d\times d}$ respectively, then by Eqs. 8 and 2.1,
@@
Chen, Nadimpalli and Yuen [Chen2023Junta, Definition 7] measured distance between unitaries by
@@
in the context of testing and learning juntas. Since
$D^{\prime}(U,V)^{2}=\frac{1}{2d}\min_{|\phi|=1}\operatorname*{tr}\mathopen{}\mathclose{{}\left((\phi U-V)\mathopen{}\mathclose{{}\left(\phi^{*}U^{\dagger}-V^{\dagger}}\right)}\right)=\min_{|\phi|=1}\mathopen{}\mathclose{{}\left(1-\frac{1}{d}\mathrm{Re}\mathopen{}\mathclose{{}\left[\phi\operatorname*{tr}\mathopen{}\mathclose{{}\left(U^{\dagger}V}\right)}\right]}\right)=1-\frac{1}{d}\mathopen{}\mathclose{{}\left|\operatorname*{tr}\mathopen{}\mathclose{{}\left(U^{\dagger}V}\right)}\right|,$
@@
i.e. $D^{\prime}$ is equivalent to $D$ (and hence to ACID distance) up to a constant factor.
@@ 4.3 Relation to distance between quantum Boolean functions
Now consider arbitrary quantum Boolean functions $F,G\in\mathbb{C}^{d\times d}$. Since $F,G$ are Hermitian it holds that $\operatorname*{tr}(FG)$ is real. Up to a ±$\pm 1$ global phase, we may further assume that $\operatorname*{tr}(FG)$ is nonnegative, and then $\Delta(F,G)=D^{\prime}(F,G)^{2}/2$ for $D^{\prime}$ defined as in Section 4.2. Since we showed in Section 4.2 that $D^{\prime}$ is proportional to ACID distance, it follows that $\Delta$ is proportional to squared ACID distance.
@@ 4.5 Relation to the induced trace norm and its average-case analogue
In Section 5 (specifically Theorem 5.5) we will prove a generalization of the fact that $\operatorname*{\mathbb{E}}{\|\mathcal{L}(\boldsymbol{\psi})\|}_{1}\leq{\|\mathcal{L}\|}_{J}$ for all superoperators $\mathcal{L}$, where $\boldsymbol{\psi}$ is Haar random. Here we give an alternate proof of this fact in the case where $\mathcal{L}$ is the difference between two unitary channels, i.e. $\mathcal{L}(X)=UXU^{\dagger}-VXV^{\dagger}$ for some unitaries $U,V\in\mathbb{C}^{d\times d}$. If $D$ denotes the average-case distance between unitaries from Section 4.2, then by Cauchy-Schwarz and Eqs. 20 and 21,
@@ 4.6 Relation to quantum fault-tolerance and experiments
We continue the discssion from the end of Section 1.4. Gilchrist, Langford and Nielsen [GLN05] proposed six properties that any distance $\Delta(\mathcal{M},\mathcal{N})$ between channels $\mathcal{M}$ and $\mathcal{N}$ should have in order to be suitable for measuring the error of a quantum computation: it should be a metric, be easy to calculate, be easy to experimentally measure, have a well-motivated physical interpretation, satisfy stability (i.e. $\Delta(\mathcal{I}\otimes\mathcal{M},\mathcal{I}\otimes\mathcal{N})=\Delta(\mathcal{M},\mathcal{N})$), and satisfy chaining (i.e. $\Delta(\mathcal{M}_{2}\mathcal{M}_{1},\mathcal{N}_{2}\mathcal{N}_{1})\leq\Delta(\mathcal{M}_{1},\mathcal{N}_{1})+\Delta(\mathcal{M}_{2},\mathcal{N}_{2})$). Kueng, Long, Doherty and Flammia [KLDF16, Eqs. 2 and 3] noted the significance of stability and chaining as well. Out of many candidate distances, Gilchrist, Langford and Nielsen [GLN05] identified four that satisfy these criteria: ACID distance (which they call Jamiołkowski process distance or $J$ distance), related distances arising from fidelity (i.e. Jamiolkowski process fidelity or $J$ fidelity), diamond distance (i.e. stabilized process distance or $S$ distance), and related distances arising from fidelity (i.e. stabilized process fidelity or $S$ fidelity). [^17] In particular, they rejected the “average-case induced trace distance” from Section 4.5 as a candidate distance [GLN05, Eq. 13], for only seeming to satisfy the metric and chaining criteria out of the six. They also gave an operational interpretation of the ACID norm as a bound on the average probability of error experienced during quantum computation of a function, or as a bound on the distance between the real and ideal joint distributions of the quantum computer in a sampling computation [GLN05, Section VI.(i)].
This suggests that one may hope to prove a fault-tolerance theorem with respect to the ACID norm. Aharonov, Kitaev and Nisan [Aharonov98, Lemma 12] listed five [^18] As well as a sixth, ${\|\mathcal{L}(\rho)\|}_{1}\leq{\|\mathcal{L}\|}_{\diamond}{\|\rho\|}_{1}$, which follows immediately from the first property and the definition of the induced trace norm. properties of the diamond norm that are used in the proof of the fault-tolerance theorem: for all superoperators $\mathcal{K},\mathcal{L}$:
@@ 5 Proof that the ACID norm is “average-case”
In this section, for an arbitrary superoperator $\mathcal{L}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ and a Haar random state $\boldsymbol{\psi}\in\mathsf{D}(d_{\mathrm{in}}\otimes d_{\mathrm{anc}})$, we give conditions under which ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\boldsymbol{\psi}\|}_{1}$ concentrates around ${\|\mathcal{L}\|}_{J}$, i.e. we prove Theorem 1.3. For technical reasons it will be convenient to refer to an unnormalized version of the maximally entangled state in this section:
@@
It will also be convenient to have a shorthand notation for the quantity ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\psi\|}_{1}$ which we are relating to ${\|\mathcal{L}\|}_{J}$. In particular, since this quantity depends only on the reduced state $\rho$ on the first register of $\psi$, it will be convenient to have a shorthand notation defined in terms of $\mathcal{L}$ and $\rho$ only. One purification [^19] I.e. a pure state whose reduced state on the first register equals $\rho$. of $\rho$ is $(\sqrt{\rho}\otimes I_{d_{\mathrm{in}}})|\Psi_{d_{\mathrm{in}}}\rangle$, as can be straightforwardly verified using the fact that the partial trace over the second register of $\Psi$ equals $I$. This motivates the following definition, which is equivalent to the trace norm of the operator defined by applying $\mathcal{L}\otimes\mathcal{I}$ to this purification of $\rho$:
@@
In particular, the $I/d$ norm is equivalent to the ACID norm (recall 1.4). In 5.2 as in the rest of this section, we denote the dimension of the input register by $d$ rather than $d_{\mathrm{in}}$ for brevity, since we will not have occasion to refer to the dimension of the output register.
@@ 5.2 Bounds on the expected $\boldsymbol{\rho}$ norm for unitarily invariant $\boldsymbol{\rho}$
$\operatorname*{\mathbb{E}}{\|\mathcal{L}\|}_{\boldsymbol{\rho}}\leq\operatorname*{tr}(\operatorname*{\mathbb{E}}\boldsymbol{\rho}\sigma)d{\|\mathcal{L}\|}_{J}={\|\mathcal{L}\|}_{J}.$
@@
where the last inequality uses linearity of expectation and the fact that $\boldsymbol{\rho}$ has unit trace.
@@
${\|\mathcal{L}(I/d)\|}_{1}={\|\mathcal{L}(\operatorname*{\mathbb{E}}\boldsymbol{\rho})\|}_{1}\leq\operatorname*{\mathbb{E}}{\|\mathcal{L}(\boldsymbol{\rho})\|}_{1},$
@@
Finally, since $d>1$ it holds that $d^{2}(2-F)-1$ is strictly positive, so we may divide both sides of the above inequality by $d^{2}(2-F)-1$, yielding
@@
$f(|\psi\rangle)-f(|\phi\rangle)={\|(\mathcal{L}\otimes\mathcal{I})\psi\|}_{1}-{\|(\mathcal{L}\otimes\mathcal{I})\phi\|}_{1}\leq{\|(\mathcal{L}\otimes\mathcal{I})(\psi-\phi)\|}_{1}\leq{\|\mathcal{L}\|}_{\diamond}{\|\psi-\phi\|}_{1}\\ \leq 2{\|\mathcal{L}\|}_{\diamond}{\||\psi\rangle-|\phi\rangle\|}_{2},$
where the last inequality is by Eq. 9. In other words $f$ is $2{\|\mathcal{L}\|}_{\diamond}$-Lipschitz, so Lemma 5.9 implies that
@@
In particular, if $m\geq\omega(\log d)$ then ${\|\mathcal{L}\|}_{\boldsymbol{\rho}}\leq O({\|\mathcal{L}\|}_{J})$ with high probability. We did not attempt to optimize the constants in Theorem 5.11.
@@
By Lemma 5.4 there exists a density matrix $\sigma\in\mathsf{D}(d)$ such that for all density matrices $\rho\in\mathsf{D}(d)$, it holds that ${\|\mathcal{L}\|}_{\rho}\leq\operatorname*{tr}(\rho\sigma)d{\|\mathcal{L}\|}_{J}$. At this point we could note that by Hölder’s inequality $\operatorname*{tr}(\boldsymbol{\rho}\sigma)\leq{\|\boldsymbol{\rho}\|}_{\infty}{\|\sigma\|}_{1}={\|\boldsymbol{\rho}\|}_{\infty}$ and bound ${\|\boldsymbol{\rho}\|}_{\infty}$ using a matrix Chernoff bound, but this is wasteful: we don’t need to bound ${\|\boldsymbol{\rho}|\phi\rangle\|}_{2}$ for every pure state $|\phi\rangle$ but only for those $|\phi\rangle$ that are eigenvectors of $\sigma$. Concretely, let $\sigma=\sum_{j=1}^{d}\lambda_{j}|\phi_{j}\rangle\!\langle\phi_{j}|$ be an eigendecomposition of $\sigma$. Then $\operatorname*{tr}(\rho\sigma)\leq\max_{j}\operatorname*{tr}(\rho\phi_{j})$ for all density matrices $\rho\in\mathsf{D}(d)$, so by a union bound [^21] A tighter but more complicated bound follows from a result of Hsu, Kakade and Zhang [HKZ12, Proposition 1].

### v2

@@ 2.3 Query models for channel testers
A (deterministic) ancilla-free, non-adaptive channel tester making $n$ queries to a channel $\mathcal{M}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ consists of the following:
@@
A (deterministic) ancilla-assisted, incoherent, non-adaptive channel tester making $n$ queries to a channel $\mathcal{M}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ consists of the following:
@@
A (deterministic) ancilla-assisted, coherent, non-adaptive channel tester making $n$ queries to a channel $\mathcal{M}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ consists of the following:
@@
A (deterministic) ancilla-assisted, coherent, adaptive channel tester making $n$ queries to a channel $\mathcal{M}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ consists of the following:
@@ 2.4 Von Neumann entropy
We will use von Neumann entropy to prove our results about tomography in Section 6.5.
@@
$\mathopen{}\mathclose{{}\left|S\mathopen{}\mathclose{{}\left(\rho^{\prime}}\right)-S\mathopen{}\mathclose{{}\left(\sigma^{\prime}}\right)}\right|\leq x\log\mathopen{}\mathclose{{}\left(d^{\prime}}\right)+1.$
@@ 3 Lower bounds for channel certification in diamond distance
It is straightforward to verify that $\boldsymbol{\mathcal{M}}$ is completely positive and trace-preserving. One may alternatively verify that $\boldsymbol{\mathcal{M}}$ is a channel by interpreting it as the following sequence of physical operations: with probability $1-\varepsilon$ apply $\mathcal{N}$, and with probability $\varepsilon$ apply the channel that first performs the PVM $(\boldsymbol{\phi},I-\boldsymbol{\phi})$ on the input state, and then outputs $\boldsymbol{\psi}$ if the measurement outcome was $\boldsymbol{\phi}$ and outputs $\mathcal{N}$ applied to the post-measurement state if the measurement outcome was $I-\boldsymbol{\phi}$. Thus $\boldsymbol{\mathcal{M}}$ behaves similarly to $\mathcal{N}$ except on inputs near $\boldsymbol{\phi}$.
@@
$\geq 2\varepsilon(1-1/d_{\mathrm{out}})$ definition of $\boldsymbol{\psi}$
$\geq\varepsilon$ $d_{\mathrm{out}}\geq 2.$
@@
where the probability is over both the choice of $\boldsymbol{\mathcal{M}}$ and the randomness of the output measurement. Recalling 2.5, write $T=(\mathcal{V}_{1},\dotsb,\mathcal{V}_{n},P)$ where $n$ is the number of queries made by $T$; our goal is to prove that $n\geq\Omega(\sqrt{d}_{\mathrm{in}}/\varepsilon)$.
@@
$\leq\varepsilon\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\mathcal{N}\mathopen{}\mathclose{{}\left((\boldsymbol{\phi}\otimes I)\xi(\boldsymbol{\phi}\otimes I)-\xi(\boldsymbol{\phi}\otimes I)-(\boldsymbol{\phi}\otimes I)\xi}\right)}\right\|_{1}+\operatorname*{tr}\mathopen{}\mathclose{{}\left((\boldsymbol{\phi}\otimes I)\xi}\right)}\right]$ triangle ineq.
@@ 4.2 Relation to average-case distance between unitaries
Throughout this subsection let $U,V\in\mathbb{C}^{d\times d}$ be arbitrary unitaries. Low [Low2009Avg, Definition 10 and Eq. 7] used the distance
$D(U,V)\coloneqq\sqrt{1-\frac{1}{d^{2}}\mathopen{}\mathclose{{}\left|\operatorname*{tr}\mathopen{}\mathclose{{}\left(U^{\dagger}V}\right)}\right|^{2}}=\frac{1}{\sqrt{2}d}{\|U\otimes U^{\dagger}-V\otimes V^{\dagger}\|}_{2}$
in the context of unitary testing and tomography. Montanaro and de Wolf [MdW13, Proposition 21] proved that
$D(U,V)=\sqrt{\frac{d+1}{4d}\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|U\boldsymbol{\psi}U^{\dagger}-V\boldsymbol{\psi}V^{\dagger}}\right\|_{1}^{2}}\right]}$ (19)
where $\boldsymbol{\psi}\in\mathsf{D}(d)$ is Haar random, giving an interpretation of $D$ as an “average-case distance”. ACID distance generalizes $D$ because if channels $\mathcal{U},\mathcal{V}\in\mathsf{C}(d,d)$ conjugate by $U$ and $V$ respectively, then by Eqs. 7 and 2.1,
@@
Zhao et al. [Zhao+23, Lemma 22 and its proof] independently observed Eqs. 19 and 20 as well. They used the distance $\sqrt{\frac{1}{4}\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|U\boldsymbol{\psi}U^{\dagger}-V\boldsymbol{\psi}V^{\dagger}}\right\|_{1}^{2}}\right]}$ in the context of unitary tomography, a quantity which is within a universal constant factor of $D(U,V)$ by Eq. 19 and the fact that $1\leq(d+1)/d\leq 2$.
Wang [Wang2011Unitary, Eq. 4] and Chen, Nadimpalli and Yuen [Chen2023Junta, Definition 7] used the distance
@@
in the context of unitary testing and tomography. Wang [Wang2011Unitary, Eq. 6] and Zhao et al. [Zhao+23, Lemma 4(1) and its proof] independently observed that
$D^{\prime}(U,V)^{2}=1-\frac{1}{d}\mathopen{}\mathclose{{}\left|\operatorname*{tr}\mathopen{}\mathclose{{}\left(U^{\dagger}V}\right)}\right|,$
@@
i.e. $D^{\prime}$ is within a constant factor of $D$ (and hence of ACID distance).
@@ 4.3 Relation to distance between quantum Boolean functions
Now consider arbitrary quantum Boolean functions $F,G\in\mathbb{C}^{d\times d}$. Since $F,G$ are Hermitian it holds that $\operatorname*{tr}(FG)$ is real. Up to a ±$\pm 1$ global phase, we may further assume that $\operatorname*{tr}(FG)$ is nonnegative, and then $\Delta(F,G)=D^{\prime}(F,G)^{2}/2$ for $D^{\prime}$ defined as in Section 4.2. Recalling that $D^{\prime}$ is proportional to ACID distance, it follows that $\Delta$ is proportional to squared ACID distance.
@@ 4.5 Relation to the induced trace norm and its average-case analogue
In Section 5 (specifically Theorem 5.5) we will prove a generalization of the fact that $\operatorname*{\mathbb{E}}{\|\mathcal{L}(\boldsymbol{\psi})\|}_{1}\leq{\|\mathcal{L}\|}_{J}$ for all superoperators $\mathcal{L}$, where $\boldsymbol{\psi}$ is Haar random. Here we give an alternate proof of this fact in the case where $\mathcal{L}$ is the difference between two unitary channels, i.e. $\mathcal{L}(X)=UXU^{\dagger}-VXV^{\dagger}$ for some unitaries $U,V\in\mathbb{C}^{d\times d}$. If $D$ denotes the average-case distance between unitaries from Section 4.2, then by Cauchy-Schwarz and Eqs. 19 and 20,
@@ 4.6 Relation to quantum fault-tolerance and experiments
We continue the discssion from the end of Section 1.4. Gilchrist, Langford and Nielsen [GLN05] proposed six properties that any distance $\Delta(\mathcal{M},\mathcal{N})$ between channels $\mathcal{M}$ and $\mathcal{N}$ should have in order to be suitable for measuring the error of a quantum computation: it should be a metric, be easy to calculate, be easy to experimentally measure, have a well-motivated physical interpretation, satisfy stability (i.e. $\Delta(\mathcal{I}\otimes\mathcal{M},\mathcal{I}\otimes\mathcal{N})=\Delta(\mathcal{M},\mathcal{N})$), and satisfy chaining (i.e. $\Delta(\mathcal{M}_{2}\mathcal{M}_{1},\mathcal{N}_{2}\mathcal{N}_{1})\leq\Delta(\mathcal{M}_{1},\mathcal{N}_{1})+\Delta(\mathcal{M}_{2},\mathcal{N}_{2})$). Kueng, Long, Doherty and Flammia [KLDF16, Eqs. 2 and 3] noted the significance of stability and chaining as well. Out of many candidate distances, Gilchrist, Langford and Nielsen [GLN05] identified four that satisfy these criteria: ACID distance (which they call Jamiołkowski process distance or $J$ distance), related distances arising from fidelity (i.e. Jamiolkowski process fidelity or $J$ fidelity), diamond distance (i.e. stabilized process distance or $S$ distance), and related distances arising from fidelity (i.e. stabilized process fidelity or $S$ fidelity). [^17] In particular, they rejected the “average-case induced trace distance” from Section 4.5 as a candidate distance [GLN05, Eq. 13], for only seeming to satisfy the metric and chaining criteria out of the six. They also gave an operational interpretation of the ACID norm as a bound on the “average probability of error experienced during quantum computation of a function, or as a bound on the distance between the real and ideal joint distributions of the quantum computer in a sampling computation” [GLN05, Section VI.(i)].
This suggests that one may hope to prove a fault-tolerance theorem with respect to the ACID norm. Aharonov, Kitaev and Nisan [Aharonov98, Lemma 12] listed five [^18] As well as a sixth, ${\|\mathcal{L}(X)\|}_{1}\leq{\|\mathcal{L}\|}_{\diamond}{\|X\|}_{1}$, which follows immediately from the first property and the definition of the induced trace norm. properties of the diamond norm that are used in the proof of the fault-tolerance theorem: for all superoperators $\mathcal{K},\mathcal{L}$:
@@ 5 Proof that the ACID norm is “average-case”
In this section we prove Theorem 1.3, i.e. we give conditions under which ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\boldsymbol{\psi}\|}_{1}$ concentrates around ${\|\mathcal{L}\|}_{J}$ for a superoperator $\mathcal{L}\in\mathsf{C}(d_{\mathrm{in}},d_{\mathrm{out}})$ and Haar random $\boldsymbol{\psi}\in\mathsf{D}(d_{\mathrm{in}}\otimes d_{\mathrm{anc}})$. For technical reasons it will be convenient to refer to an unnormalized version of the maximally entangled state in this section:
@@
It will also be convenient to have a shorthand notation for the quantity ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\psi\|}_{1}$ which we are relating to ${\|\mathcal{L}\|}_{J}$. In particular, since this quantity depends only on the reduced state $\rho$ on the first register of $\psi$, it will be convenient to have a shorthand notation in terms of $\mathcal{L}$ and $\rho$ only. One purification [^19] I.e. a pure state whose reduced state on the first register equals $\rho$. of $\rho$ is $(\sqrt{\rho}\otimes I_{d_{\mathrm{in}}})|\Psi_{d_{\mathrm{in}}}\rangle$, as can be straightforwardly verified using the fact that the partial trace over the second register of $\Psi$ equals $I$. This motivates the following definition, which is equivalent to the trace norm of the operator defined by applying $\mathcal{L}\otimes\mathcal{I}$ to this purification of $\rho$, and which generalizes the ACID norm (by taking $\rho=I/d$):
@@ 5.2 Bounds on the expected $\boldsymbol{\rho}$ norm for unitarily invariant $\boldsymbol{\rho}$
$\operatorname*{\mathbb{E}}{\|\mathcal{L}\|}_{\boldsymbol{\rho}}\leq\operatorname*{tr}(\operatorname*{\mathbb{E}}[\boldsymbol{\rho}]\sigma)d{\|\mathcal{L}\|}_{J}={\|\mathcal{L}\|}_{J}.$
@@
where the last equality uses linearity of expectation and the fact that $\boldsymbol{\rho}$ has unit trace.
@@
${\|\mathcal{L}(I/d)\|}_{1}={\|\mathcal{L}(\operatorname*{\mathbb{E}}[\boldsymbol{\rho}])\|}_{1}\leq\operatorname*{\mathbb{E}}{\|\mathcal{L}(\boldsymbol{\rho})\|}_{1},$
@@
Finally, since $d>1$ the quantity $d^{2}(2-F)-1$ is strictly positive, so we may divide both sides of the above inequality by $d^{2}(2-F)-1$, yielding
@@
$f(|\psi\rangle)-f(|\phi\rangle)$ $={\|(\mathcal{L}\otimes\mathcal{I})\psi\|}_{1}-{\|(\mathcal{L}\otimes\mathcal{I})\phi\|}_{1}$
$\leq{\|(\mathcal{L}\otimes\mathcal{I})(\psi-\phi)\|}_{1}$ triangle inequality
$\leq{\|\mathcal{L}\|}_{\diamond}{\|\psi-\phi\|}_{1}$
$\leq 2{\|\mathcal{L}\|}_{\diamond}{\||\psi\rangle-|\phi\rangle\|}_{2}$ Eq. 8.
In other words $f$ is $2{\|\mathcal{L}\|}_{\diamond}$-Lipschitz, so by Lemma 5.9
@@
We did not attempt to optimize the constants in Theorem 5.11.
@@
By Lemma 5.4 there exists a density matrix $\sigma\in\mathsf{D}(d)$ such that for all density matrices $\rho\in\mathsf{D}(d)$, it holds that ${\|\mathcal{L}\|}_{\rho}\leq\operatorname*{tr}(\rho\sigma)d{\|\mathcal{L}\|}_{J}$. At this point we could note that by Hölder’s inequality $\operatorname*{tr}(\boldsymbol{\rho}\sigma)\leq{\|\boldsymbol{\rho}\|}_{\infty}{\|\sigma\|}_{1}={\|\boldsymbol{\rho}\|}_{\infty}$ and bound ${\|\boldsymbol{\rho}\|}_{\infty}$ using a matrix Chernoff bound, but this is wasteful: we don’t need to bound ${\|\boldsymbol{\rho}|\phi\rangle\|}_{2}$ for every pure state $|\phi\rangle$, but only for those $|\phi\rangle$ that are eigenvectors of $\sigma$. Concretely, let $\sigma=\sum_{j=1}^{d}\lambda_{j}|\phi_{j}\rangle\!\langle\phi_{j}|$ be an eigendecomposition of $\sigma$. Then $\operatorname*{tr}(\rho\sigma)\leq\max_{j}\operatorname*{tr}(\rho\phi_{j})$ for all density matrices $\rho\in\mathsf{D}(d)$, so by a union bound [^21] A tighter but more complicated bound follows from a result of Hsu, Kakade and Zhang [HKZ12, Proposition 1].